# MERGE: Matching Electronic Results with Genuine Evidence for verifiable voting in person at remote locations

### v3

@@ 1 Introduction
This paper describes the MERGE protocol, which enhances the basic scheme of electronic delivery and paper returns in a way that increases speed without significantly compromising privacy or integrity. The main idea is to send an untrusted electronic record back quickly, which can be confirmed if the paper ballot arrives in time for the audit, or distrusted otherwise. This extends the time available for the paper ballot to arrive.
@@
  • The untrusted electronic record can be incorporated into preliminary results and used as the cast vote record in a Risk Limiting Audit (RLA).
@@
The paper and electronic records could be produced by scanning a hand marked paper ballot, or printing a paper ballot from a ballot marking device. [^2] Scientific opinion is divided on whether human verification of BMDs is sufficiently accurate to justify the assumption that the printout matches the voter’s intention. See [2] and [20] for examples of opposing views. At a minimum, this requires careful design to ensure voters are motivated and encouraged to verify, and there is something they can do if a misprint occurs. See section 3 for a detailed discussion of the trust assumptions.
@@ 1.1 Authentication using CAC cards
Voters at Remote Voting Centers own a smart card called a Common Access Card (CAC) card. It provides a Public Key Infrastructure (PKI) to enable secure authentication and digital signatures. Each CAC card includes a (private) signing key and a certificate linking its ID to the corresponding public key. Observers and officials in each Local Counting Center know the CAC IDs of the voters they should include, and can validate CAC certificates. Only votes with a valid CAC signature are accepted via the electronic channel.
@@ 2.1 Cryptographic components
We also use NIZK proofs to demonstrate the mixnet input $\mathcal{I}$ (a set of vote vectors) is correctly shuffled and re-encrypted to $\mathcal{O}$, without revealing any information about the permutation from $\mathcal{I}$ to $\mathcal{O}$. This proof is denoted by $\textit{ZKP}_{\textsc{Mix}}(\mathcal{I},\mathcal{O})$
@@ 2.2 Risk Limiting Audit
An RLA technique that is particularly useful in our setting is the phantoms-to-zombies approach devised by Bañuelos and Stark [3]. This is a way of dealing with ballots that appear in the manifest but cannot be found on paper, a problem that may occur frequently when MERGE ballots are delayed or dropped in the mail. The idea is to “Pretend that the audit actually finds a ballot, an evil zombie ballot that shows whatever would increase the P-value the most.” Bañuelos and Stark prove this to be conservative, in the sense that the RLA property still holds, assuming that it would have held had the correct ballot paper been located. The correct evil zombie ballot might be slightly different depending on the kind of RLA, but is generally a vote for the highest loser, or possibly an imaginary vote for all the losers, or (in the case of separate comparisons between the winner and each loser) a vote for whichever loser is being compared.
@@
An RLA functionality is monotonic if for any given setup parameters with an incorrect apparent outcome and any input sample set, an increase in the discrepancy value of any sampled ballot, while other discrepancy values remain fixed, does not change the output from ’reject’ to ’accept’.
@@ 3.1 Attacker model
The Electronic&Drop model probably corresponds most closely to intuitive assumptions about postal mail voting (whether those assumptions are valid in practice or not). Jurisdictions vary greatly in how they prevent fraudulent ballots from being accepted. Our protocol’s use of digital signature stickers makes it harder for an external attacker to generate apparently-valid fake votes, but since we assume that the attacker controls all the computers, it also controls the computers that get access to the voter’s CAC signing key. We therefore do not make it harder for this internal attacker to stuff a valid-looking paper ballot into the paper mail than it would have been with traditional postal mail. If the jurisdiction was diligently checking each voter’s handwritten signature, or if some other form of registered or secure mail is being used, our protocol does not undermine it, but nor do we add to the defences. So we simply assume that something prevents the internal attacker from adding fraudulent mail ballots.
@@ 4.1 Voting process for voter $i$
    • verify the digital signature on the sticker, and
@@ 4.2 Digital path
We explain digital processing of electronic ballots for one Local Counting Center’s votes. Every other Local Counting Center’s data is processed similarly—they do not interact because their results need to be delivered to the appropriate Local Counting Center, even for statewide contests. At the conclusion of the voting process, the following digital record containing the information of all $n$ voters becomes publicly available on the BB, where the record with index $i$ denotes the ballot from voter $i$, where $i=1\ldots{}n$.
[…]
@@
    The validity of the signature and ZKP for each electronic ballot on the BB is verified. It is also verified that no two ballots on the bulletin board share the same ciphertext.
@@ 4.3 Paper path
    Each incoming envelope’s sticker is scanned and its corresponding digital signature is verified. It is necessary to check that the data being signed matches the corresponding data on the BB for this voter and the voting computer’s signature is valid. Then, the envelope is accepted and its corresponding record on the BB is marked as received. If any of these verifications fails, the process continues but the envelope is set aside unopened in a stack called “Rejected Envelopes” for further investigations. If the signature is valid but another validly signed envelope has already been received from the same voter, it is set aside in a stack called the “Eligibility Problem” stack.
@@ 4.4.1 Determining discrepancies
    For ballots of type “duplicated”, the above 2-step procedure is performed for all $j^{\prime}$ values and the final disrepancy is set to the minimum discrepancy among different $j^{\prime}$ values for the digital ballot.
@@ 4.6 Counting the number of votes and creating a trustworthy ballot manifest
This ensures that only authorized votes are recorded on the BB. However, it does not guarantee protection against dropping ballots from the electronic record. This could be disastrous because, if the paper records are also delayed or dropped, there is no way to know how many ballots have been dropped, and consequently what the implications for the accuracy of the election result might be. It is a requirement of any RLA that there must be a record of how many ballot papers there are, derived independently of the scanners that are being audited. This assumption needs to be supported by one of the following practical procedures.
@@ 5.1 Verifiability
These statements, based on Remarks 2 and 3 respectively, imply that for any given risk limit, when the announced outcome is wrong, if the RLA functionality with any given sample set accepts the announced outcome in a newly defined world, it will accept the corresponding outcome in the present world.
@@ 5.2 Privacy
We use an instance of the privacy definition BPRIV from [8], in which the algorithms for Setup, Vote, Valid, Publish, Tally and Verify are derived from the MERGE protocol. Importantly, the Tally protocol is defined to publish both the decrypted vote tally from Step 3 and also the random permutation of individual votes from Step 5.
@@ 6 Using MERGE with other RLA styles, avoiding the publication of MERGE sub-tallies
If a Local Counting Center has only a very small number of MERGE ballots, it may not be acceptable to publish their tally in item 3. When an attacker has direct access to the arriving mail ballot, we do not have a solution to this issue. However, if we consider a remote attacker who is looking at the sub-tallies on the BB, we can address the problem by including the MERGE tally into a larger tally of ballots. In more details, the ordinary ballots from the Local Counting Center can simply be appended to the MERGE ballots and then the tallies are computed over the whole set of votes. Appendix D describes multiple variants that obviate the need to publish the sub-tally of MERGE ballots in item 3 of the MERGE digital path. Some also avoid printing individual serial numbers on ballots, relying on batch numbers that identify a group of ballots instead. Appendix D.1 describes how to incorporate MERGE into audits using VAULT [7], retaining serial numbers, mixing, and individual ballot-level comparisons. Appendix D.2 describes how to incorporate MERGE ballots into larger batches and audit them as a batch—this requires neither mixing nor serial numbers, but may result in a less efficient audit.
@@ 7 Implementation results
In this section we discuss our prototype library for the MERGE protocol. The prototype library [^12] Available at https://github.com/JohnLCaron/egk-ec-mixnet has an ElectionGuard 2.0 implementation, as well as Elliptic Curve (NIST P-256) and mixnet implementations derived from the Verificatum library [^13] See https://www.verificatum.org/, along with an option to use GMP [^14] See https://gmplib.org/ for the low-level computations, also derived from Verificatum.
@@ 9 Acknowledgements
We would like to thank Dan Wallach for valuable ideas in the early phases of this project. Thanks also to Josh Benaloh, Michael Naehrig, Olivier Pereira, John Sebes and Andrew Appel for helpful comments on earlier drafts of this paper, and to Thomas Haines for review of the mixnet code. Finally, the UI team at Rice and the development team at VotingWorks significantly influenced this protocol for the better.
@@ B.1 ElectionGuard details
To address scenarios where certain guardians may be unavailable, a robust cryptographic mechanism is employed. Guardians cryptographically share their secret keys, ensuring that a pre-defined threshold quorum guardians is required to achieve full decryption.
@@ B.2 Notations and building blocks
Similar to ElectionGuard, encryption of votes in this work is performed using a variant exponential form of the ElGamal cryptosystem. The cryptosystem parameters $(p,q,g)$ are defined as follows: $p$ is a prime number equal to $12kq+$ 1, where $q$ is also a prime number, and $g$ is a generator of the order $q$ subgroup of $\mathbb{Z}_{p}^{*}$. To generate a public-private key pair, a private key $s\in\mathbb{Z}_{q}$ is randomly selected. The corresponding public key, $K$, is then computed as $K=g^{s}\mod p$ and made public. To encrypt a vote $b$, a nonce $\xi$ is selected randomly such that $0\leq\xi<q$. Then, encryption of $b$, denoted by $\mathit{E}(b)$, is computed as
@@
B.3 Encryption scheme
The parameters of the ElGamal cryptosystem $\mathcal{G}=(p,q,g)$ are specified as follows: $p$ is a prime number given by $12kq+$ 1, where $q$ is also a prime number, and $g$ represents a generator of the order $q$ subgroup of $\mathbb{Z}_{p}^{*}$. ElGamal cryptosystem includes four main algorithms as follows:
ElGamal key generation $\mathsf{KeyGen}(\mathcal{G})$
Input: ElGamal parameters $\mathcal{G}=(p,q,g)$
Output: Public-private key pair $(K,s)$
  1. 1.
    $s\xleftarrow{\$}\mathbb{Z}^{*}_{q}$
  2. 2.
    $K=g^{s}\mod p$
  3. 3.
    Output $(K,s)$
ElGamal encryption
Input: ElGamal parameters $\mathcal{G}=(p,q,g)$, message $\sigma\in\mathbb{Z}^{*}_{q}$, public key $K$
Output: $\mathit{E}(\sigma)$
  1. 1.
    $\xi\xleftarrow{\$}\mathbb{Z}^{*}_{q}$
  2. 2.
    $\alpha=g^{\xi}\mod p,\beta=K^{\sigma}K^{\xi}\mod p$
  3. 3.
    Output $(\alpha,\beta)$
ElGamal encryption aggregation
Input: ElGamal parameters $\mathcal{G}=(p,q,g)$, encryption values $(\alpha_{i},\beta_{i})=\mathit{E}(\sigma_{i})$
Output: $\mathit{E}(\Sigma_{i}\sigma_{i}\mod q)$
  1. 1.
    $A=\prod_{i}\alpha_{i}\mod p,B=\prod_{i}\beta_{i}\mod p$
  2. 2.
    Output $(A,B)$
ElGamal decryption
Input: ElGamal parameters $\mathcal{G}=(p,q,g)$, ciphertext $C=(A,B)$, private key $s$
Output: $\mathit{D}(C;s)$
  1. 1.
    $T=\frac{B}{A^{s}}\mod p$
  2. 2.
    Output $T$
B.4 Hash computation
The function $H$ that is used in this whole section is the same as one specified in ElectionGuard. It is based on hashed message authentication code HMAC-SHA-256. The function $H$ takes two inputs, $B_{0}$ and $B_{1}$, where $B_{0}$ is 256 bit long and corresponds to the key in HMAC and $B_{1}$, which is of arbitrary length, is the actual input to the HMAC. These inputs are separated by a semicolon. If $B_{1}$ is comprised of multiple elements (e.g. $B_{1}=a||b||c$), then these elements are separated by commas( e.g. $H(B_{0};a,b,c)$).
@@ B.4 Hash computation
Input: $B_{0}$ and $B_{1}=B_{11}||B_{12}||\cdots||B_{1n}$
Output: $H(B_{0};B_{1})$
  1. 1.
[…]
B.5 Distributed ElGamal
Distributed ElGamal is a variant of the traditional ElGamal encryption scheme that involves multiple parties collaboratively generating the encryption keys and decrypting the ciphertexts.
Key generation
Principals: $n$ guardians $G_{i}$ with $i=\{1,\cdots,n\}$
Input: ElGamal parameters $\mathcal{G}=(p,q,g)$
Output: Public key shares $K_{i}$ for $i=\{1,\cdots,n\}$, aggregated public key $K$
Output ($G_{i}$): Private key share $s_{i}$
  1. 1.
    $G_{i}$: $(K_{i},s_{i})\xleftarrow[]{}\mathsf{KeyGen}(\mathcal{G})$.
  2. 2.
    $G_{i}$: Send $K_{i}$ and $\mathsf{KnowDlog}(g,K_{i})$ to $G_{j}$ for any $j\neq i$
  3. 3.
    $G_{i}$: Proceed if all proofs $\mathsf{KnowDlog}(g,K_{i})$ for $i=\{1,\cdots,n\}$ are verified. Otherwise, stop.
  4. 4.
    $G_{i}$: Output private key share $s_{i}$
  5. 5.
    All guardians: Output public key shares $K_{i}$ for $i=\{1,\cdots,n\}$ the aggregated public key $K=\prod_{i}K_{i}\mod p$
Decryption
Principals: $n$ guardians $G_{i}$ with $i=\{1,\cdots,n\}$
Public input: ElGamal parameters $\mathcal{G}=(p,q,g)$, ciphertext $C=(A,B)$, public key shares $K_{i}$ for $i=\{1,\cdots,n\}$
Private input ($G_{i}$): Private key share $s_{i}$
Output: $\mathsf{DistDec}(C,\sum_{i=1}^{n}s_{i})$
  1. 1.
    $G_{i}$: Publish share $M_{i}=A^{s_{i}}\mod p$ and proof $\mathsf{EqDlogs}(g,A,K_{i},M_{i})$.
  2. 2.
    Proceed if all proofs $\mathsf{EqDlogs}(g,A,K_{i},M_{i})$ for $i=\{1,\cdots,n\}$ are verified. Otherwise, stop.
  3. 3.
    Compute $M=\prod_{i}M_{i}\mod p$
  4. 4.
    Output $T=B/M\mod p$
Required operations per guardian: 1 exp., 1 $\mathsf{EqDlogs}$, $n-1$ $\mathsf{EqDlogs}$ verification: in total 3 exp, $2(n-1)$ multi-exp
Required operations for verification: $n$ $\mathsf{EqDlogs}$ verification: $2n$ multi-exp
Note that in the protocol presented above, anyone with access to the public input and the proofs $\mathsf{EqDlogs}(g,A,K_{i},M_{i})$ for $i=\{1,\cdots,n\}$ can verify the correctness of the distributed decryption operation by following steps 2-4 of the protocol. With the following protocol, guardians can collaboratively provide a proof of decryption correctness that can be verified by anyone who has access to the aggregated public key $K$ rather than each individual guardian’s public key share.
Proof of decryption correctness
Principals: $n$ guardians $G_{i}$ with $i=\{1,\cdots,n\}$ and the verifier $V$
Public input: ElGamal parameters $\mathcal{G}=(p,q,g)$, ciphertext $C=(A,B)$, public key shares $K_{i}$ for $i=\{1,\cdots,n\}$, aggregated public key $K$ and two deployment-dependant constants $cons_{0}$ and $cons_{1}$
Private input ($G_{i}$): Private key share $s_{i}$
Output (guardians): Plaintext $T:=\mathsf{DistDec}(C;K)$ and proof $\textit{ZKP}_{\textsc{Dec}}(T,C)$
  1. 1.
    $G_{i}$: Compute share $M_{i}=A^{s_{i}}\mod p$ and send $M_{i}$ to $G_{j}$ for any $j\neq i$
  2. 2.
    $G_{i}$: Compute $M=\prod_{j}M_{j}\mod p$, $T=\frac{B}{M}\mod p$, $u_{i}\xleftarrow{\$}\mathbb{Z}^{*}_{q}$, $a_{i}=g^{u_{i}}\mod p$ and $b_{i}=A^{u_{i}}\mod p$ and send $(a_{i},b_{i})$ to $G_{j}$ for any $j\neq i$
  3. 3.
    $G_{i}$: Compute $a=\prod_{j}a_{j}$ and $b=\prod_{j}b_{j}$ and $c=H(cons_{0};cons_{1},K,A,B,a,b,M)$
  4. 4.
    $G_{i}$: Send $v_{i}=u_{i}-cs_{i}$ to $G_{j}$ for any $j\neq i$
  5. 5.
    $G_{i}$: Compute $a^{\prime}_{j}=g^{v_{j}}K_{j}^{c}$ and $b^{\prime}_{j}=A^{v_{j}}M_{j}^{c}$ and verify if $a_{j}=a^{\prime}_{j}$ and $b_{j}=b^{\prime}_{j}$ for any $j\neq i$. Otherwise, stop.
  6. 6.
    All guardians: Compute $v=\sum_{j}v_{j}$ and send the plaintext $T$ and the proof $(c,v)$ to $V$
  7. 7.
    $V$: Compute $M=\frac{B}{T}\mod p$, $a=g^{v}K^{c}\mod p$ and $b=A^{v}M^{c}\mod p$ and verify if $v\in\mathbb{Z}_{q}$ and $c=H(cons_{0};cons_{1},K,A,B,a,b,M)$. Otherwise, reject.
Required operations per guardian: 3 exp., $2(n-1)$ multi-exp
Required operation for verification: 2 multi-exp
B.6 Zero-knowledge proofs
$\mathsf{KnowDlog}(g,K)$
Principals: The prover $P$ and the verifier $V$
Public input: ElGamal parameters $\mathcal{G}=(p,q,g)$, $K$ and two deployment-dependant constants $cons_{0}$ and $cons_{1}$
Private input ($P$): $s$ s.t. $K=g^{s}\mod p$
  1. 1.
    $P$: computes
    • $u\xleftarrow{\$}\mathbb{Z}_{q}$
    • $h=g^{u}\mod p$
    • $c=H(cons_{0};cons_{1},K,h)$
    • $v=u-cs\mod q$
  2. 2.
    $P\xrightarrow{c,v}V$
  3. 3.
    $V$: Compute $h=g^{v}K^{c}\mod p$ and then verify if $c=H(cons_{0};cons_{1},K,h)$. Otherwise, reject.
Required operations: 1 exp
Required operation for verification: 1 multi-exp
$\mathsf{EqDlogs}(x,y,X,Y)$
Principals: The prover $P$ and the verifier $V$
Public input: ElGamal parameters $\mathcal{G}=(p,q,g)$, $x$, $y$, $X$, $Y$ and two deployment-dependant constants $cons_{0}$ and $cons_{1}$
Private input ($P$): $s$ s.t. $X=x^{s}\mod p$ and $Y=y^{s}\mod p$
  1. 1.
    $P$: compute
    • $u\xleftarrow{\$}\mathbb{Z}_{q}$
    • $h=x^{u}\mod p$
    • $h^{\prime}=y^{u}\mod p$
    • $c=H(cons_{0};cons_{1},x,y,X,Y,h,h^{\prime})$
    • $v=u-cs\mod q$
  2. 2.
    $P\xrightarrow{h,h^{\prime},c,v}V$
  3. 3.
    $V$: Compute $h=x^{v}X^{c}\mod p$ and $h^{\prime}=y^{v}Y^{c}\mod p$ and verify if $c=H(cons_{0};cons_{1},x,y,X,Y,h,h^{\prime})$. Otherwise, reject.
Required operations: 2 exp
Required operation for verification: 2 multi-exp
@@ B.6 Zero-knowledge proofs
Principals: The prover $P$ and the verifier $V$
Public input: ElGamal parameters $\mathcal{G}=(p,q,g)$, ciphertext $(\alpha,\beta)$, public key $K$ and two deployment-dependant constants $cons_{0}$ and $cons_{1}$
Private input ($P$): $\xi$, $l$ s.t. $(\alpha,\beta)=(g^{\xi},K^{l}K^{\xi})$
  1. 1.
    $P$:
    • $u_{j}\xleftarrow{\$}\mathbb{Z}_{q}$ for $j=\{0\cdots,L\}$
    • $c_{j}\xleftarrow{\$}\mathbb{Z}_{q}$ for $j=\{0\cdots,L\},j\neq l$
    • $t_{j}:=(u_{j}+(l-j)c_{j})\mod q$, $(a_{j},b_{j}):=(g^{u_{j}}\mod p,K^{t_{j}}\mod p$ for $j=\{0\cdots,L\},j\neq l$,
    • $(a_{l},b_{l}):=(g^{u_{l}}\mod p,K^{u_{l}}\mod p$
    • $c=H(cons_{0};cons_{1},K,\alpha,\beta,a_{0},b_{0},a_{1},b_{1},\cdots,a_{L},b_{L})$
    • $c_{l}:=(c-\sum_{j\neq l}c_{j})\mod q$
    • $v_{j}=(u_{j}-c_{j}\xi)\mod q$
    • Send $(c_{j},v_{j})$ for $j=\{0,\cdots,L\}$ to $V$
  2. 2.
    $V$:
    1. (a)
      $(a_{j},b_{j}):=(g^{v_{j}}\alpha^{c_{j}}\mod p,K^{v_{j}-jc_{j}}\beta^{c_{j}}\mod p)$
    2. (b)
      Verify if $\sum_{j}c_{j}=H(cons_{0};cons_{1},K,\alpha,\beta,a_{0},b_{0},a_{1},b_{1},\cdots,a_{L},b_{L})$. Otherwise, reject.
Required operations: $2L$ exp
Required operation for verification: $2L$ multi-exp
@@
Principals: The prover $P$ and the verifier $V$
Public input: ElGamal parameters $\mathcal{G}=(p,q,g)$, ciphertext $(\alpha,\beta)=(g^{\xi},K^{\xi})$, public key $K$ and two deployment-dependant constants $cons_{0}$ and $cons_{1}$
Private input ($P$): $\xi$ s.t. $(\alpha,\beta)=(g^{\xi},K^{\xi})$
  1. 1.
    $P$:
    • $u_{0},u_{1},c_{1}\xleftarrow{\$}\mathbb{Z}_{q}$
    • $(a_{0},b_{0})=(g^{u_{0}}\mod p,K^{u_{0}}\mod p)$, $(a_{1},b_{1})=(g^{u_{1}}\mod p,K^{u_{1}-c_{1}}\mod p)$
    • $c=H(cons_{0};cons_{1},K,\alpha,\beta,a_{0},b_{0},a_{1},b_{1})$
    • $c_{0}=c-c_{1}$, $v_{0}=u_{0}-c_{0}\xi\mod q$, $v_{1}=u_{1}-c_{1}\xi\mod q$
    • Send $(c_{0},c_{1},v_{0},v_{1})$ to $V$
  2. 2.
    $V$
    1. (a)
      $(a_{0},b_{0}):=(g^{v_{0}}\alpha^{c_{0}}\mod p,K^{v_{0}}\beta^{c_{0}}\mod p)$
    2. (b)
      $(a_{1},b_{1}):=(g^{v_{1}}\alpha^{c_{1}}\mod p,K^{v_{1}-c_{1}}\beta^{c_{1}}\mod p)$
    3. (c)
      Verify if $c_{0}+c_{1}=H(cons_{0};cons_{1},K,\alpha,\beta,a_{0},b_{0},a_{1},b_{1})$. Otherwise, reject.
Required operations: 4 exp
Required operation for verification: 4 multi-exp

### v4

@@ 1 Introduction
This paper describes the MERGE protocol, which enhances the basic scheme of electronic delivery and paper returns in a way that increases speed without significantly compromising privacy or integrity. The main idea is to send an untrusted electronic record back quickly, which can be confirmed if the paper ballot arrives in time for the Risk Limiting Audit (RLA), or distrusted otherwise. The RLA is a post-election auditing process that uses statistical methods to ensure with high confidence that the reported election outcome is correct, by manually comparing a random sample of paper ballots against the electronic vote tallies. Since RLA normally happens a few days to a few weeks after the election day, this extends the time available for the paper ballot to arrive.
@@
  • The untrusted electronic record can be incorporated into preliminary results and used as the cast vote record in an RLA.
@@
The paper and electronic records could be produced by scanning a hand marked paper ballot, or printing a paper ballot from a ballot marking device (BMD). [^2] Scientific opinion is divided on whether human verification of BMDs is sufficiently accurate to justify the assumption that the printout matches the voter’s intention. See [2] and [20] for examples of opposing views. At a minimum, this requires careful design to ensure voters are motivated and encouraged to verify, and there is something they can do if a misprint occurs. See section 3 for a detailed discussion of the trust assumptions.
@@ 1.1 Authentication using CAC cards
Voters at Remote Voting Centers own a smart card called a Common Access Card (CAC) card. It provides a Public Key Infrastructure (PKI) to enable secure authentication and digital signatures. Each CAC card includes a (private) signing key and a certificate linking its ID to the corresponding public key. Observers and officials at each Local Counting Center know the CAC IDs and certificates of the voters to be included. This knowledge is obtained either through individual CAC voters, who must register their public keys with their local electoral authorities, or via a more centrally organized process. Only votes with a valid CAC signature are accepted via the electronic channel.
Changes to CAC cards during the protocol, such as name changes, eligibility changes or key refreshes between the time the person votes and the time the votes are tallied, are out of scope for this paper but will need to be handled with specific policies in practice.
@@ 2.1 Cryptographic components
We also use NIZK proofs to demonstrate that the mixnet input $\mathcal{I}$ (a set of vote vectors) is correctly shuffled and re-encrypted to $\mathcal{O}$, without revealing any information about the permutation from $\mathcal{I}$ to $\mathcal{O}$. This proof is denoted by $\textit{ZKP}_{\textsc{Mix}}(\mathcal{I},\mathcal{O})$
@@ 2.2 Risk Limiting Audit
An RLA technique that is particularly useful in our setting is the phantoms-to-zombies approach devised by Bañuelos and Stark [3]. This is a way of dealing with ballots that appear in the manifest but cannot be found on paper, a problem that may occur frequently when MERGE ballots are delayed or dropped in the mail. The idea is to “Pretend that the audit actually finds a ballot, an evil zombie ballot that shows whatever would increase the risk value the most.” Bañuelos and Stark prove this to be conservative, in the sense that the RLA property still holds, assuming that it would have held had the correct ballot paper been located. The correct evil zombie ballot might be slightly different depending on the kind of RLA, but is generally a vote for the highest loser, or possibly an imaginary vote for all the losers, or (in the case of separate comparisons between the winner and each loser) a vote for whichever loser is being compared.
@@
An RLA functionality is monotonic if for any given setup parameters with an incorrect apparent outcome and any input sample set, an increase in the discrepancy [^7] The difference between the voter-verified paper record and the reported electronic results for a specific ballot value of any sampled ballot, while other discrepancy values remain fixed, does not change the output from ’reject’ to ’accept’.
@@ 3.1 Attacker model
The Electronic&Drop model probably corresponds most closely to intuitive assumptions about postal mail voting (whether those assumptions are valid in practice or not). Jurisdictions vary greatly in how they prevent fraudulent ballots from being accepted. Our protocol’s use of digital signature stickers makes it harder for an external attacker to generate apparently-valid fake votes, but we assume that the attacker controls all the computers, including those to which the voter’s CAC card is attached. Consequently, the attacker can manipulate a compromised voting machine with the voter’s CAC card to sign any arbitrary message. We therefore do not make it harder for this internal attacker to stuff a valid-looking paper ballot into the paper mail than it would have been with traditional postal mail. If the jurisdiction was diligently checking each voter’s handwritten signature, or if some other form of registered or secure mail is being used, our protocol does not undermine it, but nor do we add to the defences. So we simply assume that something prevents the internal attacker from adding fraudulent mail ballots.
@@ 4.1 Voting process for voter $i$
    • verify the digital signature and the signed data on the sticker, and
@@ 4.2 Digital path
We explain digital processing of electronic ballots for one Local Counting Center’s votes. Every other Local Counting Center’s data is processed similarly—they do not interact because their results need to be delivered to the appropriate Local Counting Center, even for statewide contests.
At the conclusion of the voting process, each jurisdiction processes the votes on the BB corresponding to CAC IDs registered in that jurisdiction.
@@
    The validity of the signature and ZKP for each electronic ballot on the BB is verified.
    It is also verified that no two ballots on the bulletin board contain a common ciphertext. If this occurs, it indicates serious misbehaviour by a voting computer—the process stops.
    If any two valid ballots come from the same CAC ID, the first one is selected and the rest are withheld from any further processing. [^10] It does not matter what definition of ‘first’ is applied, as long as only one is chosen, and all $n$ selected for processing by a given jurisdiction are unambiguously marked on the BB.
@@
    The digital record on the BB now contains the information of all $n$ voters, where the record with index $i$ denotes the ballot from voter $i$, where $i=1\ldots{}n$.
[…]
  3. 3.
@@ 4.3 Paper path
    Each incoming envelope’s sticker is scanned and its corresponding digital signature is verified. It is necessary to check that the data being signed matches the corresponding data on the BB for this voter, that the voting computer’s signature is valid, and that the CAC ID is registered to vote in this jurisdiction. Then, the envelope is accepted and its corresponding record on the BB is marked as received. If any of these verifications fails, the process continues but the envelope is set aside unopened in a stack called “Rejected Envelopes” for further investigations. If the signature is valid but another validly signed envelope has already been received from the same voter, it is set aside in a stack called the “Eligibility Problem” stack.
@@ 4.4 RLA
    For ballots of type “duplicated”, the above 2-step procedure is performed for all $j^{\prime}$ values and the final discrepancy is set to the minimum discrepancy among different $j^{\prime}$ values for the digital ballot. In other words, in case of several paper ballots, one of the ballots with lowest discrepancy value is selected.
@@ 4.6 Counting the number of votes and creating a trustworthy ballot manifest
However, there is still the possibility of an attack in which a malicious voting computer colludes with some other machine that (at some point) has CAC access, in order to fabricate an apparently-valid digital vote from a voter who did not intend to vote (and did not appear at the polling place). Since this is electronically indistinguishable from valid voting, it must be defended by human procedures.
  • Officials at the Remote Voting Center must keep a count of the total number of properly completed votes.
It is a requirement of any RLA that there must be a record of how many ballot papers there are, derived independently of the scanners that are being audited. Otherwise, if the paper records are also delayed or dropped, there is no way to know how many ballots have been dropped, and consequently what the implications for the accuracy of the election result might be.
There also needs to be a way to verify that ballots have not been dropped from the electronic record. This needs to be supported by one of the following practical procedures.
@@ 5.1 Verifiability
These statements, based on Remarks 2 and 3 respectively, imply that for any given risk limit, when the announced outcome is wrong, if the RLA functionality with any given sample set accepts the announced outcome in a newly defined world, it will accept the corresponding outcome in the present world. The intuition is that inconsistencies between CVRs and paper records are handled in such a way that the overall risk value does not decrease from one scenario to the next.
@@ 5.2 Privacy
We use an instance of the privacy definition BPRIV from [8], in which the algorithms for Setup, Vote, Valid, Publish, Tally and Verify are derived from the MERGE protocol. Importantly, the Tally protocol is defined to publish both the decrypted vote tally from Step 4 and also the random permutation of individual votes from Step 6.
@@ 6 Using MERGE with other RLA styles, avoiding the publication of MERGE sub-tallies
If a Local Counting Center has only a very small number of MERGE ballots, it may not be acceptable to publish their tally in item 4. When an attacker has direct access to the arriving mail ballot, we do not have a solution to this issue. However, if we consider a remote attacker who is looking at the sub-tallies on the BB, we can address the problem by including the MERGE tally into a larger tally of ballots. In more details, the ordinary ballots from the Local Counting Center can simply be appended to the MERGE ballots and then the tallies are computed over the whole set of votes. Appendix D describes multiple variants that obviate the need to publish the sub-tally of MERGE ballots in item 4 of the MERGE digital path. Some also avoid printing individual serial numbers on ballots, relying on batch numbers that identify a group of ballots instead. Appendix D.1 describes how to incorporate MERGE into audits using VAULT [7], retaining serial numbers, mixing, and individual ballot-level comparisons. Appendix D.2 describes how to incorporate MERGE ballots into larger batches and audit them as a batch—this requires neither mixing nor serial numbers, but may result in a less efficient audit.
@@ 7 Implementation results
In this section we discuss our prototype library for the MERGE protocol. The prototype library [^14] Available at https://github.com/JohnLCaron/egk-ec-mixnet has an ElectionGuard 2.0 implementation, as well as Elliptic Curve (NIST P-256) and mixnet implementations derived from the Verificatum library [^15] See https://www.verificatum.org/, along with an option to use GMP [^16] See https://gmplib.org/ for the low-level computations, also derived from Verificatum. Note that this is a prototype library for proof-of-concept purposes, not a production-quality cryptographic library. For example, we have not (yet) implemented hash canonicalization as per the ElectionGuard specification.
@@ 9 Acknowledgements
We would like to thank Dan Wallach for valuable ideas in the early phases of this project. Thanks also to Josh Benaloh, Michael Naehrig, Olivier Pereira, John Sebes and Andrew Appel for helpful comments on earlier drafts of this paper, and to Thomas Haines for review of the mixnet code. The UI team at Rice and the development team at VotingWorks greatly improved the practicality and usability of the system. Finally, careful reviews by Trail of Bits and various anonymous reviewers significantly influenced this protocol for the better.
@@ B.1 ElectionGuard details
To address scenarios where certain guardians may be unavailable, a robust cryptographic mechanism is employed. Guardians cryptographically share their secret keys, ensuring that a pre-defined threshold quorum guardians is required to achieve full decryption. See [6] for details of the key generation protocol.
MERGE follows the ElectionGuard 2.0 specification for primitives, as described in [6]. We give a brief overview of the encryption scheme here.
@@ B.2 Notations and building blocks
In ElectionGuard, encryption of votes is performed using a variant exponential form of the ElGamal cryptosystem. The cryptosystem parameters $(p,q,g)$ are defined as follows: $p$ is a prime number equal to $12kq+$ 1, where $q$ is also a prime number, and $g$ is a generator of the order $q$ subgroup of $\mathbb{Z}_{p}^{*}$. (The ElectionGuard specification requires particular values of $p$ and $q$.) To generate a public-private key pair, a private key $s\in\mathbb{Z}_{q}$ is randomly selected. The corresponding public key, $K$, is then computed as $K=g^{s}\mod p$ and made public. To encrypt a vote $b$, a nonce $\xi$ is selected randomly such that $0\leq\xi<q$. Then, encryption of $b$, denoted by $\mathit{E}(b)$, is computed as
@@
The function $H$ that is used in this whole section is the same as one specified in ElectionGuard, which is based on HMAC-SHA-256. The function $H$ takes two inputs, $B_{0}$ and $B_{1}$, where $B_{0}$ is 256 bit long and corresponds to the key in HMAC and $B_{1}$, which is of arbitrary length, is the actual input to the HMAC. These inputs are separated by a semicolon. If $B_{1}$ is comprised of multiple elements (e.g. $B_{1}=a||b||c$), then these elements are separated by commas( e.g. $H(B_{0};a,b,c)$).
Zero knowledge proofs
We use the following ElectionGuard zero knowledge proofs, which are assumed to be paramaterized by the ElectionGuard group parameters $(p,q,g)$.
  • Distributed proof of decryption correctness, used by the guardians to decrypt output ciphertexts.
  • $\mathsf{KnowDlog}(K)$: a proof of knowledge of the discrete log of $K$ base $g$ (that is, of $x$ s.t. $K=g^{x}\bmod p$).
  • $\mathsf{EqDlogs}(x,y,X,Y)$: a proof of equality of discrete logs.